# Riemannian Geometry of Optimal Rebalancing in Dynamic Weight Automated Market Makers

### 09 Mar 2026 v2

@@
Abstract
In Temporal Function Market Making (TFMM), a dynamic weight AMM pool rebalances from initial to final holdings by creating a series of arbitrage opportunities whose total cost depends on the weight trajectory taken. We show that the per-step arbitrage loss is the KL divergence between new and old weight vectors, meaning the Fisher–Rao metric is the natural Riemannian metric on the weight simplex. The loss-minimising interpolation under the leading-order expansion of this KL cost is SLERP (Spherical Linear Interpolation) in the Hellinger coordinates $\eta_{i}=\sqrt{w_{i}}$, i.e. a geodesic on the positive orthant of the unit sphere traversed at constant speed. The SLERP midpoint equals the (AM+GM)/normalise heuristic of prior work [1], so the heuristic lies on the geodesic. This identity holds for any number of tokens and any magnitude of weight change; using this link, all dyadic points on the geodesic can be reached by recursive AM-GM bisection without trigonometric functions. SLERP’s relative sub-optimality on the full KL cost is proportional to the squared magnitude of the overall weight change and to $1/f^{2}$, where $f$ is the number of interpolation steps.
1 Introduction
Dynamic AMM pools rebalance their holdings by introducing arbitrage opportunities that disappear once reserves match the desired target. In Temporal Function Market Making (TFMM) [2], this rebalancing is driven by changing the weights of a geometric mean market maker (G3M) pool from block to block.
The pool’s asset management strategy produces new target weights at a given cadence. [^1] This can be any long-only zero-leverage strategy, for example: track a market-cap weighted index, or a trend-following strategy. Spreading a weight update over multiple intermediate steps reduces the total arbitrage cost [2], so the pool interpolates towards the target.
Prior work [1] derived the optimal weight midpoint for a weight change in the limit of small steps, and from this obtained approximately-optimal trajectories using arithmetic and geometric mean interpolation: the (AM+GM)/normalise heuristic. This heuristic captures ${\sim}95\%$ of the improvement available from full numerical optimisation, but the geometric reason for its effectiveness was left unexplained.
This paper provides that explanation. We show that:
@@ 1 Introduction
  • The loss-minimising interpolation under the leading-order expansion is SLERP (Spherical Linear Interpolation) [3] in Hellinger coordinates (Corollary 2).
  • The SLERP midpoint equals the (AM+GM)/normalise midpoint exactly, for any number of tokens (Theorem 2), explaining why the heuristic works.
  • A trig-free recursive bisection algorithm computes exact SLERP trajectories at power-of-two step counts (Corollary 3).
  • SLERP’s sub-optimality on the exact KL cost is bounded at $O(\Omega^{4}/f^{3})$ (Theorem 3).
@@
2 Background
2.1 Dynamic Weight Automated Market Makers
@@
Consider an NN-token G3M pool with weights $\boldsymbol{\mathbf{w}}=(w_{1},\ldots,w_{N})$ on the simplex ($\sum_{i}w_{i}=1$, $w_{i}>0$) and reserves $\boldsymbol{\mathbf{R}}\in\mathbb{R}^{N}_{+}$. The pool invariant is $\prod_{i}R_{i}^{w_{i}}=k$ [4, 5]. When the pool is in equilibrium, its quoted prices match market prices.
When weights change directly from $\boldsymbol{\mathbf{w}}^{\mathrm{start}}$ to $\boldsymbol{\mathbf{w}}^{\mathrm{end}}$ at constant market prices, an arbitrage opportunity is created. After arbitrage, the reserves satisfy [1]
$R_{i}^{\mathrm{end}}=R_{i}^{\mathrm{start}}\frac{w_{i}^{\mathrm{end}}}{w_{i}^{\mathrm{start}}}\prod_{j=1}^{N}\left(\frac{w_{j}^{\mathrm{start}}}{w_{j}^{\mathrm{end}}}\right)^{w_{j}^{\mathrm{end}}}.$ (1)
@@
Within each block-to-block step, market prices are modelled as constant. If, instead of updating weights directly from $\boldsymbol{\mathbf{w}}^{\mathrm{start}}$ to $\boldsymbol{\mathbf{w}}^{\mathrm{end}}$ in one step (i.e. across two blocks), the pool goes via an intermediate value $\tilde{\boldsymbol{\mathbf{w}}}$ (i.e. across three blocks), so that it is arbitraged twice, $\boldsymbol{\mathbf{w}}^{\mathrm{start}}\rightarrow\tilde{\boldsymbol{\mathbf{w}}}\rightarrow\boldsymbol{\mathbf{w}}^{\mathrm{end}}$, the total cost paid to arbitrageurs is reduced [2]. $\tilde{\boldsymbol{\mathbf{w}}}$ can take a broad range of possible values and still improve the pool’s performance. Dividing up a weight change in this way is beneficial for the same reason that splitting a large order reduces market impact: each sub-step faces a smaller price displacement (Figure 1).
Figure 1: Why weight interpolation reduces rebalancing cost. (a) For a G3M pool, rebalancing cost is quadratic in the weight change $\Delta w$, so splitting into $f$ equal sub-steps of $\Delta w/f$ reduces total cost by a factor of $f$. (b) The quadratic cost curve is flat near the origin: small weight changes cost almost nothing, while a single large change incurs a disproportionately high arb loss.
If we assume that the weight change from $\boldsymbol{\mathbf{w}}^{\mathrm{start}}\rightarrow\boldsymbol{\mathbf{w}}^{\mathrm{end}}$ is small, one can obtain that the optimal intermediate weights for the two-step process are
$\tilde{w}_{i}^{*}=\frac{w_{i}(t_{f})}{W_{0}\left(\frac{ew_{i}(t_{f})}{w_{i}(t_{0})}\right)}$ (2)
where $W_{0}(\cdot)$ is the principal branch of the Lambert W function and $e$ is Euler’s constant [1].
From bounding this result from above and below by the arithmetic and geometric means respectively, one can obtain an approximately-optimal $f$-step weight trajectory $\{\breve{\boldsymbol{\mathbf{w}}}(t_{k})\}_{k=1,...,f-1}$ from $\boldsymbol{\mathbf{w}}(t_{0})=\boldsymbol{\mathbf{w}}^{\mathrm{start}}$ to $\boldsymbol{\mathbf{w}}(t_{f})=\boldsymbol{\mathbf{w}}^{\mathrm{end}}$ as the average of the linear interpolation between these values and the geometric interpolation between these values [1]:
$w_{i}^{\mathrm{AM}}(t_{k})$ $=(1-\frac{k}{f})w_{i}(t_{0})+\frac{k}{f}w_{i}(t_{f}),$ (3)
$w_{i}^{\mathrm{GM}}(t_{k})$ $={(w_{i}(t_{0}))^{(1-\frac{k}{f})}(w_{i}(t_{f}))^{\frac{k}{f}}},$ (4)
$\breve{w}_{i}(t_{k})$ $=\frac{w_{i}^{\mathrm{AM}}(t_{k})+w_{i}^{\mathrm{GM}}(t_{k})}{\sum_{j=1}^{N}\left({w_{j}^{\mathrm{AM}}(t_{k})+w_{j}^{\mathrm{GM}}(t_{k})}\right)}.$ (5)
2.2 Information Geometry and the Fisher–Rao metric
The probability simplex $\Delta^{N-1}=\{\mathbf{p}\in\mathbb{R}^{N}_{+}:\sum_{i}p_{i}=1\}$ carries a natural Riemannian structure induced by the Kullback–Leibler (KL) divergence. For two distributions $p,q\in\Delta^{N-1}$, the KL divergence
@@ 2.2 Information Geometry and the Fisher–Rao metric
$D_{\mathrm{KL}}(p\,\|\,p+\mathrm{d}p)\;=\;\tfrac{1}{2}\sum_{i=1}^{N}\frac{(\mathrm{d}p_{i})^{2}}{p_{i}}\;+\;O(\|\mathrm{d}p\|^{3}).$ (6)
This expansion defines the Fisher–Rao metric $g$. On the ambient space $\mathbb{R}^{N}_{+}$, the metric is diagonal, $g_{ij}(p)=\delta_{ij}/p_{i}$, with the corresponding squared infinitesimal line element,
$\mathrm{d}s^{2}=\sum_{i=1}^{N}\frac{(\mathrm{d}p_{i})^{2}}{p_{i}}.$ (7)
The geometry of the simplex $(\Delta^{N-1},g)$ is that of the positive orthant of an (N1N-1)-sphere. By the theorem of Chentsov [6], this is the unique Riemannian metric (up to a scale factor) invariant under sufficient statistics, establishing it as the natural information geometry of the simplex.
3 Arbitrage loss is a KL divergence
@@
[Retention Ratio] For an NN-asset G3M pool with weights changing from $\boldsymbol{\mathbf{w}}^{\mathrm{start}}$ one block to $\boldsymbol{\mathbf{w}}^{\mathrm{end}}$ the next block, with zero fees, perfect arbitrage and, constant prices, the ratio of initial value to final value of the pool is the retention ratio
$r:=\prod_{j=1}^{N}\left(\frac{w_{j}^{\mathrm{start}}}{w_{j}^{\mathrm{end}}}\right)^{w_{j}^{\mathrm{end}}}.$ (8)
@@
$\log r=\sum_{i=1}^{N}w_{i}^{\mathrm{end}}\log\!\frac{w_{i}^{\mathrm{end}}}{w_{i}^{\mathrm{start}}}=D_{\mathrm{KL}}\!\left(\boldsymbol{\mathbf{w}}^{\mathrm{end}}\,\|\,\boldsymbol{\mathbf{w}}^{\mathrm{start}}\right).$ (9)
@@
Taking $\log$ of Eq (8) and recognising the definition of $D_{\mathrm{KL}}$. Non-negativity follows from Gibbs’ inequality. ∎
The weight vectors have the properties of distributions on the token set, and this identification holds for weight changes of any magnitude. The Fisher–Rao metric is the Hessian of the KL divergence at the diagonal [7]; its quadratic form gives the loss kernel (Eq (10) below).
@@
$\log r\;\approx\;\sum_{i=1}^{N}\frac{(\Delta w_{i})^{2}}{2\,w_{i}}.$ (10)
@@
The coordinate transformation $\eta_{i}=\sqrt{w_{i}}$ maps $\Delta^{N-1}$ isometrically (up to a factor) onto the positive orthant of the unit sphere $S^{N-1}_{+}$, under which the Fisher–Rao metric reduces to a multiple of the round metric. [^2] This map $w_{i}\mapsto\sqrt{w_{i}}$ is the Hellinger embedding [7]. Geodesics on the sphere are great circles, computable in closed form via spherical linear interpolation (SLERP) [3]. We exploit this correspondence throughout: interpolation is performed on $S^{N-1}_{+}$ and mapped back to weights by squaring.
@@
$\eta_{i}=\sqrt{w_{i}},\qquad\text{so that}\qquad w_{i}=\eta_{i}^{2},\quad\mathrm{d}w_{i}=2\eta_{i}\,\mathrm{d}\eta_{i},$ (11)
$\Rightarrow\mathrm{d}s^{2}=\sum_{i=1}^{N}\frac{4\eta_{i}^{2}\,(\mathrm{d}\eta_{i})^{2}}{\eta_{i}^{2}}=4\sum_{i=1}^{N}(\mathrm{d}\eta_{i})^{2}.$ (12)
@@
Figure 2: The Hellinger embedding maps the weight simplex $\Delta^{N-1}$ to the positive orthant of the unit sphere $S^{N-1}_{+}$ via $\eta_{i}=\sqrt{w_{i}}$. Under this isometry, the Fisher–Rao metric becomes a scaling of the standard round metric, and geodesics (shortest paths under the arbitrage cost) become great circles, computable in closed form via SLERP.
4 SLERP on the sphere under Hellinger embedding

### 31 Mar 2026 v3

@@
Abstract.
We show that when a dynamic-weight AMM rebalances by creating arbitrage opportunities, the per-step log loss is the KL divergence between successive weight vectors. The Fisher–Rao metric is therefore the natural Riemannian metric on the weight simplex. The loss-minimising interpolation under the leading-order expansion of this KL cost is SLERP (Spherical Linear Interpolation) in the Hellinger coordinates $\eta_{i}=\sqrt{w_{i}}$: a geodesic on the positive orthant of the unit sphere, traversed at constant speed. The SLERP midpoint equals the (AM+GM)/normalise heuristic of prior work [Willetts and Harrington, 2024a], so the heuristic lies on the geodesic. This identity holds for any number of tokens and any magnitude of weight change; using this link, all dyadic points on the geodesic can be reached by recursive AM-GM bisection without trigonometric functions. SLERP’s relative sub-optimality on the full KL cost is proportional to the squared magnitude of the overall weight change and to $1/f^{2}$, where $f$ is the number of interpolation steps. Under driftless GBM prices, the fractional value loss from each weight update is price-independent, and the cross term between weight and price changes telescopes, so the constant-price geometry carries over. LVR exposure introduces a finite optimal step count $f^{*}$, which lies in the perturbative regime where SLERP remains near-optimal.
1. Introduction
Dynamic AMM pools rebalance their holdings by introducing arbitrage opportunities that disappear once reserves match the desired target. In Temporal Function Market Making (TFMM) [team, 2023a], this rebalancing is driven by changing the weights of a geometric mean market maker (G3M) pool from block to block.
The pool’s asset management strategy produces new target weights at a given cadence. [^1] This can be any long-only zero-leverage strategy, for example: track a market-cap weighted index, or a trend-following strategy. Spreading a weight update over multiple intermediate steps reduces the total arbitrage cost [team, 2023a], because the cost is convex in the step size, so the pool interpolates towards the target.
Prior work [Willetts and Harrington, 2024a] derived the optimal weight midpoint in the limit of small steps, and from this obtained approximately-optimal trajectories via arithmetic and geometric mean interpolation: the (AM+GM)/normalise heuristic. This heuristic captures ${\sim}95\%$ of the improvement available from full numerical optimisation, but the geometric reason for its effectiveness was left unexplained.
This paper provides that explanation. Against a backdrop of constant prices, we show that:
@@ 1. Introduction
  • The loss-minimising interpolation under the leading-order expansion is SLERP (Spherical Linear Interpolation) [Shoemake, 1985] in Hellinger coordinates (Corollary 5).
  • The SLERP midpoint equals the (AM+GM)/normalise midpoint exactly, for any number of tokens (Theorem 6); this is why the heuristic works.
  • Exact SLERP trajectories at power-of-two step counts can be computed by trig-free recursive bisection (Corollary 7).
  • SLERP’s sub-optimality on the exact KL cost vanishes as $O(1/f^{3})$ in the step count (Theorem 9).
We then use these results to attack the case where prices are stochastic, and show that:
  • Under driftless GBM prices (martingale assumption), the retention ratio is price-independent. The cross term between weight and price changes telescopes, so the path optimisation is unchanged (§4).
  • LVR exposure introduces a finite optimal step count $f^{*}$, which lies in the perturbative regime where SLERP remains near-optimal (§5).
@@
2. Background
2.1. Dynamic Weight Automated Market Makers
@@
A block is the atomic time unit of the blockchain: all state (reserves, weights, prices) is fixed within a block and may change between consecutive blocks. [^2] On Ethereum, blocks are produced every 12 seconds; on L2 chains such as Base, every 2 seconds.
Consider an $N$-token G3M pool with weights $\boldsymbol{\mathbf{w}}=(w_{1},\ldots,w_{N})$ on the simplex ($\sum_{i}w_{i}=1$, $w_{i}>0$) and reserves $\boldsymbol{\mathbf{R}}\in\mathbb{R}^{N}_{+}$. The pool invariant is $\prod_{i}R_{i}^{w_{i}}=k$ [Martinelli and Mushegian, 2019, Evans, 2019]. When the pool is in equilibrium, its quoted prices match market prices.
When weights change directly from $\boldsymbol{\mathbf{w}}^{\mathrm{start}}$ to $\boldsymbol{\mathbf{w}}^{\mathrm{end}}$ at constant market prices, an arbitrage opportunity is created. After arbitrage, the reserves satisfy [Willetts and Harrington, 2024a]
(1) $R_{i}^{\mathrm{end}}=R_{i}^{\mathrm{start}}\frac{w_{i}^{\mathrm{end}}}{w_{i}^{\mathrm{start}}}\prod_{j=1}^{N}\left(\frac{w_{j}^{\mathrm{start}}}{w_{j}^{\mathrm{end}}}\right)^{w_{j}^{\mathrm{end}}}.$
@@
Within each block-to-block step, market prices are modelled as constant. If, instead of updating weights directly from $\boldsymbol{\mathbf{w}}^{\mathrm{start}}$ to $\boldsymbol{\mathbf{w}}^{\mathrm{end}}$ in one step (i.e. across two blocks), the pool goes via an intermediate value $\tilde{\boldsymbol{\mathbf{w}}}$ (i.e. across three blocks), so that it is arbitraged twice, $\boldsymbol{\mathbf{w}}^{\mathrm{start}}\rightarrow\tilde{\boldsymbol{\mathbf{w}}}\rightarrow\boldsymbol{\mathbf{w}}^{\mathrm{end}}$, the total cost paid to arbitrageurs is reduced [team, 2023a]. $\tilde{\boldsymbol{\mathbf{w}}}$ can take a broad range of possible values and still improve the pool’s performance. Dividing up a weight change in this way is beneficial for the same reason that splitting a large order reduces market impact [Almgren and Chriss, 2000]: each sub-step faces a smaller price displacement (Figure 1).
Figure 1. Why weight interpolation reduces rebalancing cost. (a) For a G3M pool, rebalancing cost is quadratic in the weight change $\Delta w$, so splitting into $f$ equal sub-steps of $\Delta w/f$ reduces total cost by a factor of $f$. (b) The quadratic cost curve is flat near the origin: small weight changes cost almost nothing, while a single large change incurs a disproportionately high arb loss.
Assuming the weight change is small, the optimal intermediate weights for the two-step process are
(2) $\tilde{w}_{i}^{*}=\frac{w_{i}(t_{f})}{W_{0}\left(\frac{ew_{i}(t_{f})}{w_{i}(t_{0})}\right)},$
as obtained in [Willetts and Harrington, 2024a], where $W_{0}(\cdot)$ is the principal branch of the Lambert W function and $e$ is Euler’s number.
Bounding this result from above and below by the arithmetic and geometric means respectively yields an approximately-optimal $f$-step weight trajectory $\{\breve{\boldsymbol{\mathbf{w}}}(t_{k})\}_{k=1,...,f-1}$ from $\boldsymbol{\mathbf{w}}(t_{0})=\boldsymbol{\mathbf{w}}^{\mathrm{start}}$ to $\boldsymbol{\mathbf{w}}(t_{f})=\boldsymbol{\mathbf{w}}^{\mathrm{end}}$ as the average of the linear interpolation between these values and the geometric interpolation between these values [Willetts and Harrington, 2024a]:
(3) $w_{i}^{\mathrm{AM}}(t_{k})$ $=(1-\frac{k}{f})w_{i}(t_{0})+\frac{k}{f}w_{i}(t_{f}),$
(4) $w_{i}^{\mathrm{GM}}(t_{k})$ $={(w_{i}(t_{0}))^{(1-\frac{k}{f})}(w_{i}(t_{f}))^{\frac{k}{f}}},$
(5) $\breve{w}_{i}(t_{k})$ $=\frac{w_{i}^{\mathrm{AM}}(t_{k})+w_{i}^{\mathrm{GM}}(t_{k})}{\sum_{j=1}^{N}\left({w_{j}^{\mathrm{AM}}(t_{k})+w_{j}^{\mathrm{GM}}(t_{k})}\right)}.$
2.2. Information Geometry and the Fisher–Rao metric
The weight simplex carries a natural geometric structure that, as we show in §3.2, governs the arbitrage cost.
The probability simplex $\Delta^{N-1}=\{\mathbf{p}\in\mathbb{R}^{N}_{+}:\sum_{i}p_{i}=1\}$ carries a natural Riemannian structure induced by the Kullback–Leibler (KL) divergence [Kullback and Leibler, 1951]. For two distributions $p,q\in\Delta^{N-1}$, the KL divergence
@@ 2.2. Information Geometry and the Fisher–Rao metric
(6) $D_{\mathrm{KL}}(p\,\|\,p+\mathrm{d}p)\;=\;\tfrac{1}{2}\sum_{i=1}^{N}\frac{(\mathrm{d}p_{i})^{2}}{p_{i}}\;+\;O(\|\mathrm{d}p\|^{3}).$
This expansion defines the Fisher–Rao metric $g$ [Fisher, 1925, Rao, 1945]. On the ambient space $\mathbb{R}^{N}_{+}$, the metric is diagonal, $g_{ij}(p)=\delta_{ij}/p_{i}$, with the corresponding squared infinitesimal line element,
(7) $\mathrm{d}s^{2}=\sum_{i=1}^{N}\frac{(\mathrm{d}p_{i})^{2}}{p_{i}}.$
The geometry of the simplex $(\Delta^{N-1},g)$ is that of the positive orthant of an (N1N-1)-sphere. By the theorem of Chentsov [1982], this is the unique Riemannian metric (up to a scale factor) invariant under sufficient statistics, establishing its role as the natural geometric structure.
3. Weights change with constant prices
3.1. Arbitrage loss is a KL divergence
@@
[Retention Ratio] For an $N$-asset G3M pool with weights changing from $\boldsymbol{\mathbf{w}}^{\mathrm{start}}$ one block to $\boldsymbol{\mathbf{w}}^{\mathrm{end}}$ the next block, with zero fees, perfect arbitrage, and constant prices, the ratio of final value to initial value of the pool is the retention ratio
(8) $r:=\prod_{j=1}^{N}\left(\frac{w_{j}^{\mathrm{start}}}{w_{j}^{\mathrm{end}}}\right)^{w_{j}^{\mathrm{end}}}.$
@@
(9) $\log r=\sum_{i=1}^{N}w_{i}^{\mathrm{end}}\log\!\frac{w_{i}^{\mathrm{end}}}{w_{i}^{\mathrm{start}}}=D_{\mathrm{KL}}\!\left(\boldsymbol{\mathbf{w}}^{\mathrm{end}}\,\|\,\boldsymbol{\mathbf{w}}^{\mathrm{start}}\right).$
@@
Take $\log$ of Eq (8) and recognise the definition of $D_{\mathrm{KL}}$. Note the asymmetry: $D_{\mathrm{KL}}(\boldsymbol{\mathbf{w}}^{\mathrm{end}}\|\boldsymbol{\mathbf{w}}^{\mathrm{start}})$ uses the new weights as the reference distribution, because the exponents $w_{j}^{\mathrm{end}}$ in the retention ratio determine the post-trade value allocation. Non-negativity follows from Gibbs’ inequality [Cover and Thomas, 2006]. ∎
Weight vectors are positive and sum to one, so they live on the probability simplex; Theorem 2 thus holds for weight changes of any magnitude. The Fisher–Rao metric is the Hessian of the KL divergence at the diagonal [Amari, 1985]; its quadratic form gives the loss kernel (Eq (10) below).
@@
(10) $\log r\;\approx\;\sum_{i=1}^{N}\frac{(\Delta w_{i})^{2}}{2\,w_{i}^{\mathrm{start}}}.$
@@
Proposition 0 (Lambert W midpoint as KL stationary point).
For $f=2$ (one intermediate weight vector $\boldsymbol{\mathbf{m}}$ between start and end), the unconstrained component-wise stationarity condition
yields exactly the Lambert W formula $\tilde{w}_{i}^{*}=w_{i}^{\mathrm{end}}/W_{0}(e\,w_{i}^{\mathrm{end}}/w_{i}^{\mathrm{start}})$ of Eq (2).
Proof.
Differentiating: $\partial/\partial m_{i}[m_{i}\log(m_{i}/w_{i}^{s})+w_{i}^{e}\log(w_{i}^{e}/m_{i})]=\log(m_{i}/w_{i}^{s})+1-w_{i}^{e}/m_{i}=0$, which rearranges to $m_{i}\exp(m_{i}/w_{i}^{s}-1)=w_{i}^{e}$, giving $m_{i}=w_{i}^{e}/W_{0}(e\,w_{i}^{e}/w_{i}^{s})$. ∎
The simplex constraint $\sum_{i}m_{i}=1$ is then imposed by renormalisation, which is why the formula is exact only for small weight changes: the full constrained problem couples the components through a Lagrange multiplier that has no closed-form solution [Willetts and Harrington, 2024a]. The Lambert W midpoint optimises the unconstrained component-wise KL cost, SLERP optimises the leading-order expansion of the constrained cost, and the two agree through second order (Table 1).
@@
The coordinate transformation $\eta_{i}=\sqrt{w_{i}}$ maps $\Delta^{N-1}$ isometrically (up to a factor) onto the positive orthant of the unit sphere $S^{N-1}_{+}$, under which the Fisher–Rao metric reduces to a multiple of the round metric. [^3] This map $w_{i}\mapsto\sqrt{w_{i}}$ is the Hellinger embedding [Hellinger, 1909, Amari, 1985]. Geodesics on the sphere are great circles, computable in closed form via spherical linear interpolation (SLERP) [Shoemake, 1985]. We exploit this correspondence throughout: interpolation is performed on $S^{N-1}_{+}$ and mapped back to weights by squaring.
@@
(11) $\eta_{i}=\sqrt{w_{i}},\qquad\text{so that}\qquad w_{i}=\eta_{i}^{2},\quad\mathrm{d}w_{i}=2\eta_{i}\,\mathrm{d}\eta_{i},$
(12) $\Rightarrow\mathrm{d}s^{2}=\sum_{i=1}^{N}\frac{4\eta_{i}^{2}\,(\mathrm{d}\eta_{i})^{2}}{\eta_{i}^{2}}=4\sum_{i=1}^{N}(\mathrm{d}\eta_{i})^{2}.$
@@
Figure 2. The Hellinger embedding maps the weight simplex $\Delta^{N-1}$ to the positive orthant of the unit sphere $S^{N-1}_{+}$ via $\eta_{i}=\sqrt{w_{i}}$. Under this isometry, the Fisher–Rao metric becomes a scaling of the standard round metric, and geodesics (shortest paths under the arbitrage cost) become great circles, computable in closed form via SLERP.
3.2. SLERP on the sphere under Hellinger embedding